2. What is the concentration of an HNO3 solution if 25.0 ml reacts with 38.5 ml of 0.145 M K2PO4?



1
Expert's answer
2022-03-16T05:05:05-0400

Solution:

Balanced chemical equation:

3HNO3 + K3PO4 → 3KNO3 + H3PO4

According to stoichiometry:

Moles of HNO3 / 3 = Moles of K3PO4

or:

Molarity of HNO3 × Volume of HNO3 = 3 × Molarity of K3PO4 × Volume of K3PO4

Therefore,

Molarity of HNO3 = 3 × Molarity of K3PO4 × Volume of K3PO4 / Volume of HNO3

Molarity of HNO3 = (3 × 0.145 M × 38.5 mL) / (25.0 mL) = 0.6699 M ≈ 0.670 M

Molarity of HNO3 = 0.670 M


Answer: The concentration of an HNO3 solution is 0.670 M